Calculate the pH of each of the following strong a
Answer #1

V1M1 = V2 M2

where V1 and M1 is the strength and volume of the solution initially. V2 and M2 is the strength and volume of the diluted solution.

M2 = V1m1/V2 =4.10 *0.21/49 = 0.018 M

pH = -log [H+] = -log[0.018] = 1.74

---------------------------------------------------------

Moles of H+ from 13mL 0.1M HBr = 0.1 *13/1000 = 0.0013 moles

moles of H+ from 17 mL 0.17M HCl = 0.17 *17/1000 mL = 0.00289 moles

total moles = 0.0013 moles+ 0.00289 moles = 0.00419 moles

molarity = 0.00419moles *1000/(13+17) = 0.14 M

pH = -log[H+] = 0.85